OpenStudy (annipuppi):

solve the system by substitution -4.5x-2y=-12.5 3.25x-y=-0.75

OpenStudy (cwrw238):

form the second equation y = 3.25x + 0.75 now replace the y in the first equation by 3.25x + 0.75 and solve for x finally substitute this value of x into the first equation and find y

OpenStudy (cwrw238):

* from

OpenStudy (annipuppi):

-4.5x-2(3.25x+0.75)=-12.5 -2(7.75x+0.75)=-12.5 -15.5x-1.5=-12.5 -15.5x=-11 x=0.7 is this part right or did I do something wrong?

OpenStudy (annipuppi):

@cwrw238

OpenStudy (annipuppi):

Nevermind, I figured it out. X=1 y=4